OpenStudy (help!!!!):

Factor this -64x^3+125=0

OpenStudy (campbell_st):

its the difference of 2 cubes the standard form is (a^3 - b^3) = (a - b)(a^2 + ab + b^2) so -64a^3 = -4a 125 = 5 hope it helps